What is the slope of a line apssing through the points (0,4) and (-3, -2)?
Answer by stanbon(75887)

What is the slope of a line passing through the points (0,4) and (-3, -2)?
----
slope = (4--2)/(0--3) = 6/3 = 2
----------
Cheers,
Stan H.
----------